OpenStudy (maheshmeghwal9):

The position co-ordinate of a particle that is confined to move along a straight line is given by \[\LARGE{x= 2t^3-24t+6}\] where 'x' is measured from a convenient origin & 't' is in seconds. Determine the distance travelled by the particle during the interval from "t=1 sec. to t =4 sec".

OpenStudy (maheshmeghwal9):

Its answer is 74meters. But how?

OpenStudy (maheshmeghwal9):

@mathslover @rajathsbhat Please help:)

OpenStudy (maheshmeghwal9):

I think we r asking to solve for distance not displacement so we can't do only put values

OpenStudy (anonymous):

any idea, @mathslover?

mathslover (mathslover):

no ! but i have the link of the solution can i give it ?

OpenStudy (maheshmeghwal9):

ya of course:)

mathslover (mathslover):

so here i go with the link :

mathslover (mathslover):

is this better one ?

mathslover (mathslover):

@rajathsbhat wht do u think ? is that right ? acc. to me it seems to be

OpenStudy (maheshmeghwal9):

@mathslover I think one point is to be noted that in the last part u have found displacement but we need distance & one more think at t=2 seconds v=0 u will see it

OpenStudy (maheshmeghwal9):

one more thing*

OpenStudy (anonymous):

*trying hard to figure out what it all means* *haven't slept in 23 hours....* --.--

mathslover (mathslover):

k np @rajathsbhat u can take rest now .. @maheshmeghwal9 let me think abt this deeper :)

OpenStudy (maheshmeghwal9):

ok:)

OpenStudy (maheshmeghwal9):

@rajathsbhat u must sleep now:)

OpenStudy (anonymous):

thx, @mathslover. Appreciate it :)

mathslover (mathslover):

ur welcome

mathslover (mathslover):

i am just totally confused :(

OpenStudy (maheshmeghwal9):

ya i too .

OpenStudy (maheshmeghwal9):

so anybody here Please help:)

OpenStudy (anonymous):

answer is 74....@maheshmeghwal9 said it at the beginning.

OpenStudy (anonymous):

got it! here's what i did: first differentiate the given exp to get v(t) = 6t^2 - 24 now integrate this graph twice - first from 1 to 2, then from 2 to 4. answer to first will come out as -10 and the second will give 64. distance is total area under the graph. so ignore the -ve sign and add the 2 values. so 10+64 = 74m

OpenStudy (anonymous):

its answer should be 74 units. The 54 units is the displacement while 74 units is the distance.

OpenStudy (anonymous):

Answer explained by @Vaidehi09 is well explained.

OpenStudy (anonymous):

yup. disp = -10+64 = 54m while dist = 10+64 = 74m

OpenStudy (anonymous):

@ramkrishna thank u :)

OpenStudy (maheshmeghwal9):

sorry i was offline for sometime but i gt ur solution @Vaidehi09 thanx a lot:)

OpenStudy (anonymous):

sure, no prob!